What is the maximum mass of S8 that can be produced by combining 77.0 g of each reactant?

8SO2+16H2S⟶3S8+16H2O
mass of S8
Answer #1

The given reaction is

8SO2 + 16H2S  \rightarrow 3S8 + 16H2O

Lets calculate the moles of reactants

Molar mass of SO2 = 64.06 g/mol

Molar mass of H2S = 34.08 g/mol

using formula Moles = mass / molar mass

Moles of SO2 = 77.0 g / 64.06 g/mol = 1.2020 mol SO2

Moles of H2S = 77.0 g/ 34.08 g/mol = 2.2594 mol H2S

Now lets see moles of H2S reacts with moles of SO2

then ,  2.2594 mol H2S would react completely with 2.2594 mol ( 3 mol S8 / 16 mol H2S ) = 0.42 mol SO2

but as per above calculation there is more SO2 present

So SO2 is in excess and H2S is limiting reactant

So

maximum mass of S8 = 2.2594 mol H2S \times (3 mol S8 / 16 mol H2S) \times molar mass of S8

= 2.2594 mol H2S \times (3 mol S8 / 16 mol H2S) \times 256.52 g/mol

= 108.67 g S8

So our answer is 108.67 g S8